# A Matter of Annotation: An Empirical Study on In Situ and Self-Recall Activity Annotations from Wearable Sensors

### v2

@@ 4.1 Missing Annotations and Consistency Across Methods
Missing Annotations and the consistency of labels set over one week varied greatly depending on the study participant. However, tendencies concerning specific methods are observable.
@@ 4.2 Visual Time-Series Analysis
The four bars that are visible above the accelerometer data are the labels set by the participants for every layer. The order is from bottom to top: in situ button, in situ app. pure self-recall, and time-series recall. Examples of labels that differ regarding the applied labeling method are marked with red boxes. The x-axis of every subplot represents roughly 8-9 hours of data. Most of the day was not labeled and is therefore categorized as void. However, such long periods often contain shorter periods of other activities, like walking. This makes it difficult to define a distinguishable void-class, which results in false positive classifications of non-void samples. Figure 5 visually shows that each participant labels his or her data very subjectively. The long green-labeled periods of participant 74e4 represent the class desk_work. The only other participant that used this label is 90a4. Since each of the study participants works in an office environment and thus conclusively works at a desk, we can assume that the same class is classified as void for all other study participants. This intra-class and inter-participant discrepancy becomes a problem whenever a model is trained that is supposed to generalize across individuals. To reduce these side effects and focus on the experiment itself, we decided to evaluate personalized models that take weekly data from participants into account.
The in situ button annotation is empty for 5 participants: eed7, 36fd, 74e4, 90a4 and d8f2. Labels are only partially set or missing entirely for this annotation method and we therefore assume that participants tend to forget to press the button on the smartwatch. Both tables, 1 and 2, support this assumption, as this labeling method shows a high percentage of missing annotations as well as a low Cohen score of 0.36% (week 1) and 0.39% (week 2). The pure self-recall method , visible on the 2nd upper layer, is often misaligned compared to the in situ methods as well as the time-series recall method . Participants tend to round up or down the start- and stop-time in steps of 5 or 10 minutes. For example, the annotations in Figure 5 given by the subjects 2b88, 834b or f30d, show such incorrectly annotated data. The pink color represents the class walking. With a closer look at the corresponding time-series data, one can see that the in situ button annotation (bottom layer) and time-series recall annotation (top layer) belongs to the typical periodic pattern of walking than the period labeled by pure self-recall.
A consistent reliable performance in all labeling methods can only be observed at the participants 4531 and fc25. Other participants like eed7, 36fd, 74e4 or a506 are very precise in their annotations across methods, but are missing at least one layer of labels. The complete collection of visualizations is available in our dataset repository444https://doi.org/10.5281/zenodo.7654684
@@ 4.3 Effects on Classification
The results of our deep learning evaluation555Detailed results for every participant included in our deep learning evaluation can be accessed online on the Weights & Biases platform: https://tinyurl.com/4vxvfaed. suggest that the annotation method chosen can have a crucial impact on the classification ability of a trained deep learning model. Depending on the chosen methodology, the average F1-Score results differ by up to 8%, as depicted in Figure 6.
In the first week, the in situ methodologies, button and app , generally perform better than the pure self-recall diary .
@@
Study participants mostly correctly estimated the duration of an activity, but tended to round up or down the start and end times. The in situ methods are up to 8% better than the pure self-recall, although the amount of annotated data available, due to missing annotations, is significantly lower than for other methods. Although, we work with a dataset recorded in-the-wild, the deep learning results generally show a high F1-Score. This is untypical for such datasets but can be explained by the fact that the majority of the daily data are assigned to the void class. This leaves proportionally only a few samples that are crucial for determining the classification performance.
Even though the number of available annotations that have been labeled by the study participants using the time-series recall method is significantly higher with 92.33%, the average F1-Score is 1.1% lower (89.00%) than the results reached with the App Assisted method (90.1%). To understand this result it is crucial to look at Table 3 in detail and take meta-information about the participants into account. The participants mostly used their diary as a mnemonic aid for the graphical annotation method and tried to identify the corresponding periods in the acceleration data. The results of subjects 2b88, a506 and eed7 show that the performance of the classifier could be increased with graphical assistance. However, the F1-Score of 2b88 is 0.01% below the F1-Score of the in situ app assisted annotation method . These subjects have in common that they are already trained in interpreting acceleration data due to their prior knowledge and thus assign samples to specific classes more precisely.
Subjects fc25, 4531, and 834b, on the other hand, do not have prior knowledge. Apart from subject 834b, the deep learning results show that presenting a visualization to an untrained participant rather harmed than helped the classifier. If one looks at the visualizations of day 1 & 6, week 2 of fc25, see Figure 5 and 7, the labels set by the subject with the help of the graphical interface, it is comprehensible that this study participant tended to be rather confused by the graphical representation and therefore labeled the data incorrectly.
@@ 5 Discussion
Through our investigation of the consistency of annotations between methodologies, we were able to show that participants in our study seem to prefer to write an activity diary (pure self-recall method ). This finding corresponds to what Vaizmann et al. [23] already points out. However, this method has the disadvantage that it can be imprecise, which is evident in the visualization of the data and annotations. Similarly, the activity diary methodology performed the least reliably among all methodologies, which has been confirmed by the deep learning model. Since the deep learning results using the in situ app annotations are almost similar to the results given by the time-series recall , even though the number of labeled samples is lower, it raises the question if a smaller set of high-quality annotations is more valuable for a classifier than a larger set of annotated data that comes with imprecise labels. This could mean that in future works we can reduce the amount of necessary training samples drastically if a certain annotation quality can be assured. However, this needs to be confirmed by further investigations.
@@
Apart from trying to solve annotation difficulties during the annotation phase itself, we can also partially counter wrong or noisy classified data by using machine learning techniques like Bootstrapping, see Miu et al. [24] or using a loss function that specifically tries to counteract this problem, such as [5, 25]. By using Bootstrapping, the machine or deep learning classifier is initially trained by a small subset of high-confident labels and further improved by using additional data. However, this technique comes with the trade-off that whenever wrong-labeled data is introduced as training data, the error will get propagated into the model. An effect that sooner or later occurs as long as the annotation methodologies themselves are not further researched. Other machine learning techniques that can work with noisy labels, see [26], are already successfully tested for Computer Vision problems and can, in theory, be adopted for Human Activity Recognition. However, earlier research has shown that not every technique that is applicable in other fields is also applicable to sensor-based data [16].
@@ 5.1 Discussing different Annotation Biases
We need to take into account that several biases could have been introduced due to the chosen annotation guidelines and tools. For example, the usage of in situ annotation methods during the day can have a positive effect on the self-recall capabilities of a participant at the end of the day. However, the comparison of consistencies across methods does not confirm that this effect indeed occurred. Every study participant showed an almost complete overall profile of self-recall annotations, even though the person has not used or has incomplete in-situ annotations, see Figure 4. However, deeper investigations are needed to be able to understand such effects better.
@@ 6 Conclusions
To guarantee high-quality annotations in future studies, especially in an uncontrolled real-world environment where a video recording is not available for the subsequent acquisition of ground truth, further research in this field is necessary. The methodologies used for annotating activity data need to be challenged and further developed.
@@
Lessons Learned. During this study, we gained insights about the effects of different annotation methods on the reliability and consistency of annotations and finally on the classifier itself, but also about training deep learning models on data recorded in-the-wild. In this chapter, we would like to share these insights to help other researchers perform their experiments more successfully. With regards to Table 4, we can narrow down specific study setups that either benefit more from self-recall or in situ annotation methods. As part of our annotation guidelines, we allowed our study participants to name their activities as they wished. Therefore, we were forced to simplify certain activities. To be able to create a real-world dataset that contains complex classes or even classes that consist of several subclasses, more elaborated annotation methods and tools must be developed. We believe that with the currently available resources, the hurdle lies very high for such datasets to be annotated accurately.
@@
Declarations
Conflict of Interest
On behalf of all authors, the corresponding author states that there is no conflict of interest.
@@
The studies involving human participants were reviewed and approved by the Ethics Committee of the University of Siegen (ethics vote #ER_12_2019). The participants provided their written informed consent to participate in this study and were briefed prior to the data capture about the study goals.
Availability of data and materials
The datasets generated during and/or analyzed during the current study are available at the Zenodo repository: https://doi.org/10.5281/zenodo.7654684.
Code availability
All code is made publicly available on our GitHub repository https://github.com/ahoelzemann/annotationMatters.
@@
Authors’ contributions

### v3

@@ 5.2 Missing Annotations and Consistency Across Methods
Missing Annotations and the consistency of labels set over the course of one week varied greatly depending on the study participant. However, tendencies with regard to specific methods are observable.
We computed missing annotations by merging all available annotations from the various methods used (button, app, diary, GUI) into an artificial global ground truth. We then compared each individual annotation layer against this consolidated ground truth to identify any missing annotations, leveraging the collective information as a reference point.
@@ 5.3 Visual Time-Series Analysis
The four bars that are visible above the accelerometer data are the labels set by the participants for every layer. The order is from bottom to top: in situ button, in situ app. pure self-recall, and time-series recall. Examples of labels that differ with regard to the applied labeling method are marked with red boxes. The x-axis of every subplot represents roughly 8-9 hours of data. Most of the day was not labeled and is therefore categorized as void. However, such long periods often contain shorter periods of other activities, like walking. This makes it difficult to define a distinguishable void-class, which results in false positive classifications of non-void samples. Figure 6 visually shows that each participant labels his or her data very subjectively. The long green-labeled periods of participant 74e4 represent the class desk_work. The only other participant that used this label is 90a4. Since each of the study participants works in an office environment and thus conclusively works at a desk, we can assume that the same class is classified as void for all other study participants. This intra-class and inter-participant discrepancy becomes a problem whenever a model is trained that is supposed to generalize across individuals. To reduce these side effects and focus on the experiment itself, we decided to evaluate personalized models that take weekly data from participants into account.
The in situ button annotation is empty for 5 participants: eed7, 36fd, 74e4, 90a4 and d8f2. Labels are only partially set or missing entirely for this annotation method and we therefore assume that participants tend to forget to press the button on the smartwatch. Both tables, 2 and 3, support this assumption, as this labeling method shows a high percentage of missing annotations as well as a low Cohen score of 0.36% (week 1) and 0.39% (week 2). The pure self-recall method , visible on the 2nd upper layer, is often misaligned compared to the in situ methods as well as the time-series recall method . Participants tend to round up or down the start- and stop-time in steps of 5 or 10 minutes. For example, the annotations in Figure 6 given by the subjects 2b88, 834b or f30d, show such incorrectly annotated data. The pink color represents the class walking. With a closer look at the corresponding time-series data, one can see that the in situ button annotation (bottom layer) and time-series recall annotation (top layer) belongs to the typical periodic pattern of walking than the period labeled by pure self-recall.
A consistent reliable performance in all labeling methods can only be observed at the participants 4531 and fc25. Other participants like eed7, 36fd, 74e4 or a506 are very precise in their annotations across methods, but are missing at least one layer of labels. The complete collection of visualizations is available in our dataset repository777https://doi.org/10.5281/zenodo.7654684
@@ 5.4 Effects on Classification
The results of our deep learning evaluation888Detailed results for every participant included in our deep learning evaluation can be accessed online on the Weights & Biases platform: https://tinyurl.com/4vxvfaed. suggest that the annotation method chosen can have a crucial impact on the classification ability of a trained deep learning model. Depending on the chosen methodology, the average F1-Score results differ by up to 8%, as depicted in Figure 7.
In the first week, the in situ methodologies, button and app , generally perform better than the pure self-recall diary . Study participants mostly correctly estimated the duration of an activity, but tended to round up or down the start and end times. The in situ methods are up to 8% better than the pure self-recall, although the amount of annotated data available, due to missing annotations, is significantly lower than for other methods. Although, we work with a dataset recorded in-the-wild, the deep learning results generally show a high F1-Score. This is untypical for such datasets but can be explained by the fact that the majority of the daily data are assigned to the void class. This leaves proportionally only a few samples that are crucial for determining the classification performance.
Even though the number of available annotations that have been labeled by the study participants using the time-series recall method is significantly higher with 92.33%, the average F1-Score is 1.1% lower (89.00%) than the results reached with the App Assisted method (90.1%). To understand this result it is crucial to look at Table 4 in detail and take meta-information about the participants into account.
@@
The participants mostly used their diary as a mnemonic aid for the graphical annotation method and tried to identify the corresponding periods in the acceleration data. The results of subjects 2b88, a506 and eed7 show that the performance of the classifier could be increased with graphical assistance. However, the F1-Score of 2b88 is 0.01% below the F1-Score of the in situ app assisted annotation method . These subjects have in common that they are already trained in interpreting acceleration data due to their prior knowledge and thus assign samples to specific classes more precisely.
Subjects fc25, 4531, and 834b, on the other hand, do not have prior knowledge. Apart from subject 834b, the deep learning results show that presenting a visualization to an untrained participant rather harmed than helped the classifier. If one looks at the visualizations of day 1 & 6, week 2 of fc25, see Figure 6 and 8, the labels set by the subject with the help of the graphical interface, it is comprehensible that this study participant tended to be rather confused by the graphical representation and therefore labeled the data incorrectly.
@@ 6 Discussion
Through our investigation of the consistency of annotations between methodologies, we were able to show that participants in our study seem to prefer to write an activity diary (pure self-recall method ). This finding corresponds to what [13] already points out. However, this method has the disadvantage that it can be imprecise, which is evident in the visualization of the data and annotations. Similarly, the activity diary methodology performed the least reliably among all methodologies, which has been confirmed by the deep learning model. Since the deep learning results using the in situ app annotations are almost similar to the results given by the time-series recall , even though the number of labeled samples is lower, it raises the question if a smaller set of high-quality annotations is more valuable for a classifier than a larger set of annotated data that comes with imprecise labels. This could mean that in future works we can reduce the amount of necessary training samples drastically if a certain annotation quality can be assured. However, this needs to be confirmed by further investigations.
@@
Apart from trying to solve annotation difficulties during the annotation phase itself, we can also partially counter wrong or noisy classified data by using machine learning techniques like Bootstrapping, see [39] or using a loss function that specifically tries to counteract this problem, such as [5, 51]. By using Bootstrapping, the machine or deep learning classifier is initially trained by a small subset of high-confident labels and further improved by using additional data. However, this technique comes with the trade-off that whenever wrong-labeled data is introduced as training data, the error will get propagated into the model. This is an effect that sooner or later occurs as long as the annotation methodologies themselves are not further researched. Other machine learning techniques that can work with noisy labels, see [52], are already successfully tested for Computer Vision problems and can, in theory, be adopted for Human Activity Recognition. However, earlier research has shown that not every technique that is applicable in other fields is also applicable to sensor-based data. [45].
Cause of Missing Annotations: We believe that specific activities are more likely to be forgotten during the labeling process than others. These activities are generally more spontaneous and require less dedicated preparation time. Examples might include classes like laying, sitting, walking, bus_driving, car_driving, eating, or desk_work. In contrast, other classes like shopping, yoga, playing_games, badminton, cooking, or horse_riding are often time-intensive, physically or mentally demanding, and frequently planned in advance or even take place at dedicated locations. Therefore, it is likely that participants find these activities easier to recall and label accurately. Obtaining separate annotations for each activity through distinct and dedicated annotation processes would have yielded valuable insights; however, this approach was deemed unfeasible for the participants involved in our study. The immense time commitment and laborious efforts required from our participants to annotate each activity individually would have imposed an unreasonable burden, rendering such a comprehensive annotation strategy impractical within the constraints of our study.
@@ 6.1 Discussing different Annotation Biases
Directly quantifying the perceived workload of subjective tasks like data labeling is a complex challenge. This difficulty stems from several factors. Firstly, individual differences in mental stamina and task perception mean what one person finds laborious, another might find manageable [53]. Secondly, memory biases can lead to under- or overestimates of effort depending on the emotional context of the task or the participant’s current state [54]. Social desirability bias can also come into play, with participants potentially downplaying their workload to appear competent or exaggerating it to justify breaks [55]. Therefore, accurately quantifying the workload associated with each of the four labeling methods presents a significant challenge. While surveys, like the NASA TLX [56], asking participants about perceived effort hold some value, these results are inherently subjective and can be heavily influenced by individual experiences and biases. While aiming for a fully objective measure of workload is desirable, it might require collecting more personal data from participants. This additional data could include details like preferred wearable devices (e.g., smartwatches), smartphone usage patterns, or individual memory recall capabilities or the emotional state of a participant [57]. While recording and quantifying this type of personal data would have provided valuable insights, it would have also significantly increased the workload placed on participants. This additional workload fell outside the scope of the current study, which prioritized collecting data through the four predefined methods.
However, we need to acknowledge that several biases could have been introduced due to the chosen annotation guidelines and tools. For example, the usage of in situ annotation methods during the day can have a positive effect on the self-recall capabilities of a participant at the end of the day. The comparison of consistencies across methods does not confirm that this effect indeed occurred. Every study participant showed an almost complete overall profile of self-recall annotations, even though the person has not used or has incomplete in situ annotations, see Figure 5. However, deeper investigations are needed to be able to understand such effects better.
@@
A Parallel annotation bias can arise in two scenarios: when multiple annotators independently label the same data, or when a single annotator uses multiple labeling methods for the same data, where the application of one method influences the subsequent labeling decisions made with other methods. There are three main ways this bias can manifest:
1.
Anchoring bias [59]: The initial labeling method might act as an anchor, subtly influencing the annotator’s decisions when using subsequent methods, even if their initial assessment might differ.
2.
Confirmation bias [60]: The annotator might subconsciously favor interpretations that align with labels generated from previous methods, overlooking alternative possibilities.
3.
Method bias [61]: Certain methods might inherently be easier or more difficult to use for specific types of data, potentially leading to systematic inconsistencies across the labeled data.
The presence of parallel annotation bias in this context suggests that the annotations might not be entirely independent between methods, potentially impacting the overall quality of the data. Anchoring and confirmation bias can lead to a lack of diversity in annotations and potentially perpetuate errors. Method bias can introduce inconsistencies that complicate data analysis. We recognize the possibility of parallel annotation bias in our dataset, where applying one labeling method might influence subsequent methods used by the same participant. However, prioritizing participant engagement, we opted for a parallel approach. This decision ensured the workload remained manageable and prevented participant dropout from the study.
@@ 7 Conclusions
High-quality annotations are crucial for accurate activity recognition, especially in uncontrolled real-world settings where video recordings are unavailable for ground truth verification. To address this challenge, further research on activity data annotation methodologies is necessary. These methodologies should empower annotators to label data in a way that comprehensively captures the subtleties of everyday life. The annotations must not only be extensive but also complete and coherent, ensuring a consistent and well-defined understanding for the AI model to learn from. Furthermore, leveraging learning methodologies like Weakly Supervised Learning methods, exemplified by works such as [62] and [63], can potentially utilize datasets like ours. However, a more comprehensive evaluation is needed to determine their suitability for real-world application.
@@
Lesson Learned. During this study, we gained insights about the effects of different annotation methods on the reliability and consistency of annotations and finally on the classifier itself, but also about training deep learning models on data recorded in-the-wild. In this chapter, we would like to share these insights to help other researchers perform their experiments more successfully. With regards to Table 5, we are able to narrow down specific study setups that either benefit more from self-recall or in situ annotation methods.
1.
Due to the good acceptance and the low workload for study participants we can recommend a self-recall method for studies where label precision is not the highest priority and rough estimations of activities are sufficient.
2.
According to our study, we can increase the label precision of the self-recall method with additional software that visualizes the raw data, e.g. [9]. We recommend considering the implementation of such a module and providing this software to participants together with an introduction on how to interpret sensor data. According to [13] the self-recall method can also be effectively improved by introducing server guesses of activities or visually organizing the day chronologically.
3.
In situ annotations result generally in more precise labels. However, the label process is more labor intensive than a self-recall method, since it can take a lot of time and often includes many steps to set the label. We argue, that smaller studies with participants who agree with performing such laborious work can benefit from this method. Such a system needs to be implemented carefully and with a holistic concept in order to not be seen as a burden by the participants [64].
4.
Annotating data with commercial apps, like Strava, is negligible due to data and privacy concerns.
5.
In situ annotation can have the same benefits as an app solution. However, only if researchers have access to the programming interface of the recording device and can implement additional features that help participants not forget to set a label.
As part of our annotation guidelines, we allowed our study participants to name their activities as they wished. Therefore, we were forced to simplify certain activities. To be able to create a real-world dataset that contains complex classes or even classes that consist of several subclasses, more elaborated annotation methods and tools must be developed. We believe that with the currently available resources, the hurdle lies very high for such datasets to be annotated accurately.
@@
Conflict of Interest Statement
The authors declare that the research was conducted in the absence of any commercial or financial relationships that could be construed as a potential conflict of interest.
Author Contributions
@@
The studies involving human participants were reviewed and approved by the Ethics Committee of the University of Siegen (ethics vote #ER_12_2019). The participants provided their written informed consent to participate in this study and were briefed before the data capture about the study goals.
@@
Supplemental Data
All code is made publicly available on our GitHub repository https://github.com/ahoelzemann/annotationMatters. Furthermore, a summary of all deep learning experiments can be accessed on the Weights and Biases https://tinyurl.com/4vxvfaed.
Data Availability Statement
The datasets generated for this study can be found in the Zenodo Repository https://doi.org/10.5281/zenodo.7654684.